OpenStudy (epicteatime):

why is sqrt(2x+3)=x equal to x=3

OpenStudy (epicteatime):

\[\sqrt{2x+3}=x\]

OpenStudy (epicteatime):

@iambatman

OpenStudy (epicteatime):

love that name XD

OpenStudy (gorv):

square both side

OpenStudy (epicteatime):

yep i know that

OpenStudy (epicteatime):

but i don't know what i do afterwards

OpenStudy (epicteatime):

do i just subtract 3?

OpenStudy (gorv):

root will cancel due to square

OpenStudy (epicteatime):

i know that

OpenStudy (gorv):

2x+3=x^2

OpenStudy (epicteatime):

yep

OpenStudy (epicteatime):

i also know that XD

OpenStudy (gorv):

now subtract 3 from that

OpenStudy (epicteatime):

yep i said that

OpenStudy (epicteatime):

then do i move 2x over there as well?

OpenStudy (epicteatime):

but then how do i get x=3

OpenStudy (gorv):

2x=x^2-3 now subtract 2x from both side 0=x^2-2x-3

OpenStudy (epicteatime):

huh

OpenStudy (epicteatime):

then?

OpenStudy (gorv):

now what we need is to factorise it

OpenStudy (epicteatime):

ohhh

OpenStudy (gorv):

x^2-3x+x-3=0

OpenStudy (epicteatime):

wait what

OpenStudy (gorv):

can u do further ??? and tell me the ans

OpenStudy (epicteatime):

WHOAAAAAaaa

OpenStudy (epicteatime):

where did the extra x come from

OpenStudy (gorv):

its done to factor

OpenStudy (epicteatime):

:|

OpenStudy (epicteatime):

thats my clueless face

OpenStudy (gorv):

we have to break -2x such that sum = -2 and product =-3

OpenStudy (gorv):

u know how to factorise

OpenStudy (epicteatime):

yes, but i still don't understand where x came from

OpenStudy (epicteatime):

or is mandatory? like is it supposed to be there?

OpenStudy (epicteatime):

or is there a place where you drag it from?

OpenStudy (gorv):

leave it and you factorise first by your way..and forget abt that x

OpenStudy (gorv):

x^-2x-3=0

OpenStudy (gorv):

do its factor

OpenStudy (epicteatime):

BUT

OpenStudy (epicteatime):

ummm

OpenStudy (epicteatime):

uhh

OpenStudy (epicteatime):

huh

OpenStudy (epicteatime):

ok i give up

OpenStudy (epicteatime):

please explain your ways master

OpenStudy (gorv):

lolll not master ....u know wat we do to factor ???

OpenStudy (epicteatime):

see thats the problem

OpenStudy (epicteatime):

i completely just brain farted

OpenStudy (epicteatime):

i don't remember XD

OpenStudy (gorv):

x^2-2x-3=0 we factor the term with x

OpenStudy (epicteatime):

but there is no other x

OpenStudy (loser66):

The simplest answer for "why" is : because when x =3, \(\sqrt {2x+3}= \sqrt {2*3+3}=\sqrt9=3\)

OpenStudy (gorv):

we factor 2 such way that sum of its factor =-2

OpenStudy (epicteatime):

whaaat

OpenStudy (epicteatime):

huh

OpenStudy (epicteatime):

BUT I NEED TO KNOW WHY THERe'S A X THERE

OpenStudy (gorv):

well @Loser66 there is no whyyy....u have to go by proper method

OpenStudy (epicteatime):

yeah, technically.

OpenStudy (loser66):

hahaha... the question is why, right?

OpenStudy (epicteatime):

I'm the visual type, so i HAVE to know why theres a x there

OpenStudy (gorv):

we factor the term with x such that ....sum of its factor =-2 and product = constant term which is =-3

OpenStudy (gorv):

so we use -2=-3+1

OpenStudy (loser66):

ok, if you need prove, here is first off, you have to have 2x+3 >0,

OpenStudy (gorv):

sum=-3+1=-2 product=(-3)*1=-3

OpenStudy (epicteatime):

yeah...

OpenStudy (gorv):

x^2+(-3+1)x-3=0 x62-3x+x-3=0

OpenStudy (loser66):

That helps you reject x =-3 after all

OpenStudy (gorv):

sorry that is x^2 not x62

OpenStudy (epicteatime):

lol

OpenStudy (gorv):

now you got ??

OpenStudy (epicteatime):

WHERE DID X62 COME FROM!!!<<DWJFKHRI"W

OpenStudy (epicteatime):

somewhat :D

OpenStudy (epicteatime):

OH I REMEMBERED FROM MY BRAIN FART

OpenStudy (epicteatime):

RIGHTTT

OpenStudy (gorv):

man thatzz what i wrote it is not x62 it is x^2

OpenStudy (epicteatime):

you have to find the pair that multiplies to 3, but adds up to -2

OpenStudy (gorv):

just go through carefuly wat i posted

OpenStudy (epicteatime):

RIGHTTTT

OpenStudy (epicteatime):

OH MY BRAIN FART JUST GOT REMOVED

OpenStudy (gorv):

multiply to -3 and add up to -2

OpenStudy (epicteatime):

OH PHEWWW

OpenStudy (epicteatime):

yeah that

OpenStudy (gorv):

lolll

OpenStudy (epicteatime):

well thanks :D

OpenStudy (epicteatime):

have a nice day or night XD

OpenStudy (gorv):

wlcm:)

OpenStudy (gorv):

hmmm night :P

OpenStudy (epicteatime):

:D